### v1

# Enhancing Sensitivity for Di-Higgs Boson Searches Using Anomaly Detection and Supervised Machine Learning Techniques

Abstract
This paper explores different strategies for enhancing sensitivity to new heavy resonances that decay into two or more Higgs bosons. This is achieved using two neural network architectures: an unsupervised autoencoder for anomaly detection and a supervised classifier. The autoencoder is trained on a small fraction of Standard Model (SM) Monte Carlo simulated events to calculate the loss distribution for input events, aiding in determining the extent to which events can be considered anomalous. The supervised classifier uses the same inputs but is trained on events simulated using both beyond Standard Model (BSM) and SM processes. By applying selection cuts to the output scores, we compare the sensitivities of the two approaches.
11footnotetext: Preprint: HEP-ANL-195813
1 Introduction
The most common approach for searches for new physics is to design an event classifier using neural networks or boosted decision trees. Such machine learning (ML) tools use events generated by Monte Carlo models, representing both Beyond Standard Model (BSM) and Standard Model (SM) processes.
Reliance on BSM physics simulations can be avoided when creating neural networks that “remember” the primary kinematic characteristics of collisions events which are predominantly governed by SM processes. If such networks are trained on a small fraction of well-understood data, or SM Monte Carlo and applied to data, the events which are recognized by such networks can be removed, leaving only those that cannot be fully reconstructed. In this case, no prior knowledge of BSM models is required, making this selection method more agnostic to potential new physics. An overview of anomaly detection methods is provided in [1]. The first LHC article utilizing anomaly detection to identify collision events with five types of reconstructed objects (jets, $b$-jets, electrons, muons and photons) that are most likely to contain BSM scenarios was recently published by ATLAS [2].
This article explores different strategies for enhancing sensitivity to new heavy resonances decaying to two or more Higgs bosons. This was done using two neural network architectures: an unsupervised autoencoder for anomaly detection (AD) and a supervised classifier (SC). The autoencoder is trained on a small fraction of SM Monte Carlo events to calculate the loss distribution for input events, helping to determine the degree to which the events can be considered anomalous. The supervised classifier uses similar inputs, but is trained on BSM and SM events. The performance of the SC is tested on events generated from a BSM model from which it has seen no training data. This is then compared with the performance of the AD model on those same BSM events, allowing for a comparison of the robustness of the two techniques.
2 Monte Carlo Simulations
For the demonstration of the event classifier and anomaly detection techniques, several Monte Carlo event samples are used as outlined below:
2.1 Signal models
The current analysis uses two exotic processes with a heavy scalar $X$ decaying into other scalars. The first process is:
$X\rightarrow HH\rightarrow VVb\bar{b}\rightarrow llb\bar{b}+A,$ (2.1)
where $H$ is the SM Higgs boson. The last part of this process indicates the decay mode, where one Higgs boson decays to vector bosons ($V$ = $Z,W$), while the second Higgs boson decays to two $b$-initiated jets (i.e. $b\bar{b}$). It is assumed that the two leptons ($ll$) in the final state are expected to originate from either the $V=Z$ boson ($Z\rightarrow l^{+}l^{-})$ or from semileptonic decays of two $V=W$ bosons. No requirements were set on the charge of the leptons. The symbol $A$ in Eq. 2.1 indicates other decay products of the vector bosons, which could be neutrinos or extra jets. They will not be used in this analysis. The considered Higgs-decay channel may not be the most competitive compared to other decay modes (such as $\gamma\gamma$); however, it provides a rich final state for studying ML techniques with complex event topologies.
The event samples with the $X\rightarrow HH$ process were created with Pythia8 version 8.307 [3] using $pp$ collisions at $\sqrt{s}=13.6$ TeV. The generated masses for the heavy scalar $X$ were set to 0.5, 0.7, 1.0, 1.5, and 2 TeV. The cross sections were set to the model double Higgs models as predicted by the MadGraph5_aMC@NLO [4] at next-to-leading order in QCD [5]. A sample of 0.5 million events was created by allowing all Higgs decays. The generated events were preselected with at least two leptons at transverse momenta of $p_{T}>14$ GeV.
In this analysis, the "signal" region represents the invariant mass of two $b-$ associated jets, which will be denoted by $M_{bb}$. This invariant mass should exhibit a sharp peak near the Higgs mass of 125 GeV.
The second signal sample was created from the model [6]:
$X\rightarrow SH\rightarrow HHH\rightarrow llb\bar{b}+A$ (2.2)
where a new scalar $S$ decays to two other SM Higgs bosons. This process was also generated with Pythia8 using $pp$ collisions at $\sqrt{s}=13.6$ TeV. This particular decay chain was recently studied in [7]. As before, the symbol $A$ indicates other decay products of $HHH$, which are not considered in this study. The generated masses for the heavy scalar $X$ were set to the same values as for the process in Eq. 2.1. The mass of the scalar $S$ was set to 280 GeV, which is slightly above the mass threshold of the $HH$ system. This selection of the mass provides the largest cross section for this channel. Then the generated events were preselected with at least two leptons at $p_{T}>14$ GeV.
In the case of the $X\to HH$ process, the invariant mass $M_{bb}$ for the process $X\to SH$ is expected to peak around the mass of the Higgs boson. This invariant mass will represent the signal region used to estimate the number of signal events from the scalar $X$, if only the first pair of jets associated with $b$-quarks is considered.
2.2 Background sample
The main SM background for the $HH$ and $HS$ processes is events with two leptons from top-quark production. Double ($t\bar{t}$) and single ($t$) top production processes were created with Pythia8. All top decays were enabled, but events were selected with at least two leptons at $p_{T}>14$ GeV. The second largest background is $W/Z$+jet production. It was also generated using Pythia8. Both SM processes were scaled to their corresponding cross sections according to the SM.
2.3 Event processing for ML inputs
Following event generation with the requirement of two leptons with $p_{T}>14$ GeV, all analysis objects were derived from truth-level particles. The jets were constructed with the anti-$k_{T}$ algorithm [8] as implemented in the FastJet package [9] with a distance parameter of $R=0.4$, which is commonly used in the ATLAS experiment. By default, the minimum transverse energy of all jets was $20$ GeV in the pseudorapidity range of $|\eta|<2.5$. Leptons were required to be isolated using a cone of size $0.2$ in the azimuthal angle and pseudorapidity defined around the true direction of the lepton. All energies of particles inside this cone were summed. A lepton is considered isolated if it carries more than $90\%$ of the cone energy. The SM background processes require simulations of misidentification rates for muons and electrons ("fake rates"). We used a misidentification rate of 0.1% for muons and 1% for electrons. This was implemented by assigning a probability of $10^{-3}$ ($10^{-2}$) for a jet to be identified as a muon (electron) using a random number generator.
The signal region of both processes is based on the measured $bbll$ final state. As the final step, all events were selected by requiring two $b$-associated jets with $p_{T}>20$ GeV and two leptons above $15$ GeV.
(a) $X\rightarrow HH$
(b) $X\rightarrow SH$
Figure 1: The $M_{bb}$ invariant mass for $X\rightarrow HH$ and $X\rightarrow SH$ processes before applying ML algorithms. The cross section for the latter process was set to that of $X\rightarrow HH$. Different shades of the gray color show the BSM signals with different masses of $X$, i.e. 0.5, 0.7, 1, 1.5 and 2 TeV. The larger the mass of $X$, the smaller area under the shaded histograms is expected.
Figure 1 shows the signal region with the invariant mass of two $b-$jets for the models $HH$ and $SH$, together with the expected SM backgrounds. The predictions for the number of signal events were obtained for the 140 fb-1 of an integrated luminosity. The number of signal events for the model $SH$ was set to that of the model $HH$ in order to simplify the comparisons for the significance calculation, which will be discussed later. As expected, both BSM models show a peak near the Higgs mass of 125 GeV. The signal from the $SH$ model is somewhat broader, due to a larger combinatorial background arising from a misassignment of $b-$jets from the Higgs decays. The signal events are difficult to observe because of the overwhelming contributions from the SM background near the 125 GeV region.
The number of $W/Z$+jet events in the signal region is roughly an order of magnitude smaller than the number from the top processes. Furthermore, it has been observed that $W/Z$+jet events are generally much easier to separate from signal events compared to top process events. For these reasons and for simplicity, the SC and AD models are trained using only top process events as background.
After all objects (jets, $b-$jets, leptons and photons) are available, a C++ program is used to transform kinematic features of each event to the rapidity–mass matrix (RMM) which is provided as an input for ML [10]. The RMM is a square matrix that includes reconstructed final states of jets, $b$-jets, muons, electrons, photons, and $E_{T}^{miss}$, where $E_{T}^{miss}$ is a single object, followed by $10$ ($b$-)jets and $5$ electrons, muons and photons each, in descending order of transverse energy for each particle type. By construction, all elements of the RMM are defined to be between 0 and 1, and most variables are Lorentz-invariant under boosts along the longitudinal axis. To reduce biases in the shapes of the jet + $Y$ invariant mass spectra, the nine invariant mass variables are excluded from the RMM. The resulting input dimension is $128736^{2}-9=$ 1287. The RMM matrix is then flattened into a one-dimensional input vector before being fed into the AE.
The RMM was first proposed for backpropagation-based event classification algorithms [10, 11]. Later, it was successfully demonstrated that the same inputs are appropriate for anomaly detection using autoencoders [12, 2, 13]. Therefore, the RMM inputs are expected to perform equally well for both supervised classifiers and unsupervised autoencoders, which is crucial for comparing these two techniques that adopt the same inputs.
2.4 Unsupervised autoencoder
The autoencoder is implemented using TensorFlow [14]. It comprises two sections, an encoder and a decoder. The encoder compresses the input to a latent dimensional space, whereas the decoder takes the data in the latent layer and decompresses it back to its original size. The network architecture for the encoder contains two hidden layers, with 800 and 400 neurons, respectively, and a latent layer of 200 neurons. The decoder reverses the structure of the encoder, using 400 and 800 neurons for the two hidden layers, and 1287 neurons for the output layer. The Leaky ReLU [15] activation function is used for all hidden and output layers. This architecture of autoencoder was found to be the most optimal for performing unsupervised learning [2, 13] using the RMM inputs. The performance can be tested using the ADfilter tool [16].
The autoencoder was trained by minimizing the MSE error using a batch size of 10. With a patience value of 10, the typical number of epochs was 200 for the final training.
2.5 Supervised classifier
The supervised classifier uses the same RMM input as the unsupervised autoencoder and was also created with TensorFlow [14]. The classifier repeats the architecture of the encoding part of the autoencoder, having 2 hidden layers with 800 and 400 neurons each and the ReLU activation function between layers. The output layer has a single neuron that was used to classify the events.
During training, the output layer was set to 0 for SM events and 1 for BSM events. The same number of SM and BSM events were used, dividing both samples into training and validation sets with a 70/30 ratio. The mass of the scalar $X$ was not included in the inputs, allowing the network to be blind to the physics parameters of the BSM model.
3 Results
Figures 2 and  3 illustrate the results of a supervised classifier and anomaly detection algorithm applied to the processes $X\to HH$ and $X\to SH$, comparing different mass hypotheses for the resonance $X$ and standard model top-quark production generated with PYTHIA8.
Figure 2 displays the score distributions for $X\to HH$ (left) and $X\to SH$ (right), where the x-axis represents the classifier’s confidence in identifying an event as a signal. The different colored lines represent various signal masses (500 GeV to 2000 GeV), while the dotted black line corresponds to the Standard Model (SM) background. The red vertical line indicates the classification threshold (0.5), where events to the right are identified as signal-like, demonstrating separation, though variations exist between $X\to HH$ and $X\to SH$.
(a) $X\rightarrow HH$
(b) $X\rightarrow SH$
Figure 2: Scores from the supervised classifier for the processes $X\rightarrow HH$ and $X\rightarrow SH$. The vertical red lines show the threshold values used to select signal events.
[…]
Figures  4 and  5 show the distributions for $X\to HH$ and $X\to SH$ after applying ML using the SC and AD, respectively. In the $X\to HH$ case, the SC accepts a slightly larger number of backgrounds events than the AD. For smaller $X$ mass points, the SC accepts a significantly larger number of signal events than the AD. The difference in signal acceptance between the models gradually decreases until the highest mass point 2 TeV, where they accept an almost identical number of events. The $X\to SH$ case is largely similar, though the difference between the SC and AD is smaller for both signal and background.
(a) $X\rightarrow HH$
(b) $X\rightarrow SH$
Figure 3: Loss scores from the anomaly detection for the processes $X\rightarrow HH$ and $X\rightarrow SH$. The vertical red lines show the threshold values used to select anomalous events using the $3\sigma$ rule based on the background events (dashed black lines).
(a) $X\to HH$ after SC
(b) $X\to HH$ after AD
Figure 4: Distributions for $X\rightarrow HH$ after applying the SC (a) and AD (b) selections.
Figure 6 shows that the significance is enhanced using the SC and AD for both $X\to HH$ and $X\to SH$ channels. For $X\to HH$ channel, the supervised event classifier is more effective at lower masses, and for higher masses, significance is enhanced using AD as well. For $X\to SH$ channel, at higher masses, AD outperforms the SC. It is expected that two b-jets from a single Higgs boson are boosted at large masses of $X$. This may reduce the efficiency of reconstruction of $M_{bb}$, but this effect should be common for both SC and AD reconstructions, and thus should not affect our conclusion about the comparison of these two ML approaches.
(a) $X\to SH$ after SC
(b) $X\to SH$ after AD
Figure 5: Distributions for $X\rightarrow SH$, with $S\rightarrow HH$ after applying the SC (a) and AD (b) selections.
(a) $X\to HH$
(b) $X\to SH$
Figure 6: Comparison of the significance values for (a) $X\rightarrow HH$; (b) $X\rightarrow SH$, with $S\rightarrow HH$ channels.
4 Conclusion
In summary, we tested two fundamentally different ML approaches: one, a supervised classifier that requires both SM background and BSM signal events as input, and the other, an autoencoder-based anomaly detection approach that only requires SM events for training. The benchmarks were based on models with $X\rightarrow HH$ and $X\rightarrow SH$ boson decays, with the final state involving two leptons and two $b$-jets. The same input variables were used in both ML approaches, and the architecture of the hidden layers was identical. Therefore, these studies provide a compelling comparison of two distinct approaches that use the same inputs.
We concluded that both ML techniques were able to improve the significance of signal observations. The event classifier outperforms the autoencoder for the BSM events used in training, especially for low masses of the scalar $X$. This is unsurprising for the $HH$ signals, which were used to train the SC. However, it was unexpected to observe that the classifier works quite well even for the $SH$ signal events, which were not used in training. Its performance is comparable to that of the autoencoder. The latter starts to slightly outperform the supervised classifier at larger masses of $X$. The good agreement between the SC and AD for the $SH$ channel is likely due to the fact that the final state of these BSM models is relatively similar. It is expected that the autoencoder may start to outperform the SC for events that have a substantial difference in the final state, such as the presence of additional identified particles that are not part of the Higgs decays. The anomaly-detection technique has not been used in di-Higgs searches. However, it offers a model-agnostic approach that generalizes more effectively across multiple BSM scenarios compared to a supervised classifier in di-Higgs analyses targeting high-mass resonances.
The Monte Carlo data and the analysis code for this paper are available from [17] and [18].
5 Acknowledgments
We would like to thank R. Zhang, D. Chen, I. Low for the discussion of the anomaly detection technique and for providing the Madgraph model cxSM_VLF_EFT for the $HHH$ signals.
The submitted manuscript has been created by UChicago Argonne, LLC, Operator of Argonne National Laboratory (“Argonne”). Argonne, a U.S. Department of Energy Office of Science laboratory, is operated under Contract No. DE-AC02-06CH11357. The U.S. Government retains for itself, and others acting on its behalf, a paid-up nonexclusive, irrevocable worldwide license in said article to reproduce, prepare derivative works, distribute copies to the public, and perform publicly and display publicly, by or on behalf of the Government. The Department of Energy will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan. http://energy.gov/downloads/doe-public-access-plan. Argonne National Laboratory’s work was funded by the U.S. Department of Energy, Office of High Energy Physics (DOE OHEP) under contract DE-AC02-06CH11357. The Askaryan Calorimeter Experiment was supported by the US DOE OHEP under Award Numbers DE-SC0009937, DE-SC0010504, and DE-AC02-76SF0051. WI is supported by DE-SC0017647. We gratefully acknowledge the computing resources provided by the Laboratory Computing Resource Center at Argonne National Laboratory.